# Thibaud Cambronne12, Samuel Bobick13, Wente Zeng4, Scott Moura12

### 23 Mar 2026 v3

@@ II Benchmark Optimization Problem
The benchmark solution is a conventional non-anticipatory optimization that involves solving (1) directly. We assume that users choose one of the 3 charging options based solely on the price menu. We denote $\text{Pr}\left(m\mid z\right)$ to be the probability that new user $n$ chooses option $m$ when presented with the set of prices $z$. We model these probabilities with a two-step discrete choice model, described in Appendix A. Due to the non-convexity of the discrete choice model (A.5)–(A.9), optimization problem (1) is difficult to solve numerically. As such, we find an approximate solution to (1) by searching over a grid of candidate prices $z_{\text{candidate}}=(z_{\text{sch}},z_{\text{reg}})$. At each point in the grid, we solve $\min_{p}\mathsf{E}[{f(p,m)\mid z_{\text{candidate}}}]$, which is a linear program.
@@ III-C Model Predictive Control
Forecasting models 2) and 3) use features such as recent historical power profiles, planned power profiles from the last optimization, number of sessions, and encodings of time of day, week, year, and workday. Model hyperparameters are tuned with a train-validation-test split of empirical station usage data. More details on feature engineering and forecast implementation can be found in Appendix B.
@@ IV-A Simulation Overview
We perform a Monte Carlo simulation [^1] Simulation code available at https://github.com/samuelBobick/StationLevelPowerForecasting. to quantitatively validate the performance of the control algorithms presented in the previous section. We replay all EV charging sessions sessions on the Smart Learning Pilot for EV Charging Stations (SlrpEV) platform from 2023, a total of 2274 charging sessions [Ozturk2025]. We assume a demand charge of \$20/kW [mclaren2017identifying], representing 24% of simulated total station operator costs under the benchmark algorithm. We use the Pacific Gas & Electric Commercial Business Electric Vehicle June 2023 rates for the TOU tariff structure [pge_tariff]. Note that the computational complexity of the algorithms proposed in Section III is not prohibitive; all algorithms run within in a fraction of a second on a standard laptop. Further details on the simulation setup can be found in Appendix C.
@@ V Conclusions and Limitations
Appendix A Behavioral Model Formulation
@@
Appendix B Time Series Forecasting Details
@@
Appendix C Simulation Setup

### 11 Apr 2026 v4

@@ Abstract
Nomenclature
  • Indices/Sets
  • $\tau/\mathcal{T}$
    Time index of the monthly billing cycle.
  • $n$
    New user whose arrival triggers current optimization.
  • $i/\mathcal{I}$
    Set of all active users connected to the charging station, including user $n$, $\mathcal{I}=\mathcal{A}_{\text{sch}}\cup\mathcal{A}_{\text{reg}}\cup\{n\}$.
  • $m$
    Charging service choice selected by user $n$. $m\in\{\textit{REGULAR}{},\textit{SCHEDULED}{},\textit{LEAVE}{}\}$.
  • $i,j/\mathcal{A}_{m}$
    Set of existing users with service option $m$.
  • $T^{\text{start}}$
    Time index of current optimization, i.e. when user $n$ arrives at the station.
  • $T_{i}^{\text{end}}$
    The time index when user $i$ is no longer charging. This corresponds to the departure time indicated by a SCHEDULED user or the time index when a REGULAR user’s battery is full.
  • $T^{\text{end}}$
    Time index of latest departure across all users $i\in\mathcal{I}$. $T^{\text{end}}=\max_{i\in\mathcal{I}}T_{i}^{\text{end}}$.
  • Parameters
  • $\Delta t$
    Time step of the system, in hours.
  • $p^{\text{max}}$
    Maximum power of a charger, in kW.
  • $\eta$
    Charger efficiency.
  • $c_{\tau}$
    Utility rate for electricity at time index $\tau$, in \$/kWh.
  • $c_{d}$
    Utility rate for demand charge, in \$/kW.
  • ${\text{E}}_{i}^{\text{req}}$
    Required energy of user $i$, in kWh.
  • Variables
  • $p$
    Power profiles of all users in $\mathcal{I}$ to be optimized in current optimization, in kW.
  • $p^{\text{sch}}_{i,\tau}$
    Charging power for SCHEDULED user $i$ at time $\tau$ in the current optimization, in kW.
  • $e_{i,\tau}$
    Energy level for user $i$ at time index $\tau$, in kWh.
  • $G_{\tau}$
    Station charging load at time index $\tau$, in kW.
  • $\hat{G}_{\tau\mid m}$
    The forecasted station charging load at time index $\tau$, assuming new user $n$ chooses charging option $m$, in kW.
  • $D_{\tau}$
    Station peak power up to time $\tau$, in kW.
  • $\zeta_{i}$
    Fixed charging price for existing user $i$ at the time index of the current optimization, in \$/kWh.
  • $z^{m}$
    Charging tariff for new user $n$ who selects option $m$, in \$/kWh.
  • $z$
    Set of prices $(z^{\text{sch}},z^{\text{reg}})$ to be presented to new user $n$, in \$/kWh.
@@ II Benchmark Optimization Problem
The benchmark solution is a conventional non-anticipatory optimization that involves solving (1) directly. We assume that users choose one of the 3 charging options based solely on the price menu. We denote $\text{Pr}\left(m\mid z\right)$ to be the probability that new user $n$ chooses option $m$ when presented with the set of prices $z$. We model these probabilities with a two-step discrete choice model, described in Appendix 0.A. Due to the non-convexity of the discrete choice model (A.5)–(A.9), optimization problem (1) is difficult to solve numerically. As such, we find an approximate solution to (1) by searching over a grid of candidate prices $z_{\text{candidate}}=(z_{\text{sch}},z_{\text{reg}})$. At each point in the grid, we solve $\min_{p}\mathsf{E}[{f(p,m)\mid z_{\text{candidate}}}]$, which is a linear program.
@@ III-C Model Predictive Control
Forecasting models 2) and 3) use features such as recent historical power profiles, planned power profiles from the last optimization, number of sessions, and encodings of time of day, week, year, and workday. Model hyperparameters are tuned with a train-validation-test split of empirical station usage data. More details on feature engineering and forecast implementation can be found in Appendix 0.B.
@@ IV-A Simulation Overview
We perform a Monte Carlo simulation [^1] Simulation code available at https://github.com/samuelBobick/StationLevelPowerForecasting. to quantitatively validate the performance of the control algorithms presented in the previous section. We replay all EV charging sessions sessions on the Smart Learning Pilot for EV Charging Stations (SlrpEV) platform from 2023, a total of 2274 charging sessions [Ozturk2025]. We assume a demand charge of \$20/kW [mclaren2017identifying], representing 24% of simulated total station operator costs under the benchmark algorithm. We use the Pacific Gas & Electric Commercial Business Electric Vehicle June 2023 rates for the TOU tariff structure [pge_tariff]. Note that the computational complexity of the algorithms proposed in Section III is not prohibitive; all algorithms run within in a fraction of a second on a standard laptop. Further details on the simulation setup can be found in Appendix 0.C.
@@ V Conclusions and Limitations
Appendix 0.A Behavioral Model Formulation
@@
Appendix 0.B Time Series Forecasting Details
@@
Appendix 0.C Simulation Setup